According to the following reaction, how many grams of bromine monochloride will be formed upon the complete reaction of 26.6 grams of chlorine gas with excess bromine?
1
Expert's answer
2016-10-19T13:53:09-0400
The equation for this reaction is:
Br2 + Cl2 → 2BrCl
The number of moles of chlorine gas equals:
ν(Cl2) = m(Cl2)/Mr(Cl2) = 26.6 g / (2×35.5 g mol-1) = 0.3746 mol
According to the equation the amount of formed BrCl is in two times more than that for Cl2.
Thus, ν(BrCl) = 2ν(Cl2) = 0.7493 mol
And its mass is:
m(BrCl) = ν(BrCl)×Mr(BrCl) = 0.7493 mol × 115 g mol-1 = 86.17 g